THE MILLIMETER/SUBMILLIMETER SPECTRUM OF AlSH (X̃^{1}A′): FUTHER INVESTIGATION OF METAL HYDROSULFIDE STRUCTURES

\underline{A. J. JANCZYK} and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721.

The pure rotational spectrum of AlSH has been recorded using millimeter/submillimeter wave direct absorption techniques. The molecule was produced by reacting aluminum vapor with H_{2}S in the presence of a d.c. discharge. Transitions ranging from J = 23 → 24 to J = 30 → 31 in the region 319 - 418 GHz were measured. The spectra exhibited K_{a} ladder structure for K_{a} = 0 to K_{a} = 6 for the species, which is consistent with the molecule being an asymmetric top. The higher asymmetry components (K_{a} = 5,6) are, however, perturbed. Rotational constants as well as an r_{o} structure have been determined. The study shows that the AlSH molecule is bent and therefore different from linear AlOH.